Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2015-03-11

Observer : Morgan Fraser
Support team : Southampton (Mat Smith and others)
TAT : Joe Anderson, Cosimo Inserra

Start of night: clear sky, but with some haze towards the horizon, also clouds behind mountaint to N-E,  5 m/s wind, 50% humidity
Moon from 02.50 onwards at 65% illumination
End of night:
Seeing: DIMM was only working intermittently, but from this an acquisition images, it was probably <1.2" for most of the night

I concentrated on EFOSC priority followup before the moon rose, then switched to SOFI to do a couple of additional critical SOFI requests which turned up. End of the night I switched back to EFOSC, as there were no SOFI targets at suitable RAs. Only one classifcation tonight.


Afternoon calibrations:    SOFI arcs and specflats

23.12    Sunset

23.25    BR skyflats

23.33    Focus

23.42    Spectrophotometric standard LTT745a

23.57    PSNJ07361576-6930230
             1x1500s Gr#11,
1x1500s Gr#16 follow-up spectra
            
From very rough photometry on the acqusition image, it looks like the SN is
             ~0.5 mag brighter than last night, supporting a young age

01.02    Re-did focus (temperature increased quite suddenly)

01.07    ASSASN-14ha
             Attempted to take followup spectra, but SN not visible on acquisition image.
             Suggest deep imaging instead - can try again tomorrow if PI wants.

01.12    PSNJ10544216-2104138
             2x1800s Gr#13 follow-up. Almost identical to 05hk.

02.17    PS15ae
             1x1800s Gr#11, 1x1800s Gr#16 follow-up spectra

03.34    Switching to SOFI
             TO tried to focus, PSF was very distorted so attempted image analysis

03.50    PSNJ07361576-6930230
             JHK phot, BG+RG spectra. Source is detected in all three filters, spectra have good
             S/N in BG, not so good in RG.
           
05.12    HIP035096 Telluric standard for PSNJ07361576-6930230

05.17    PS15br
             JHK imaging

06.34    SN2015D
             JHK imaging

07.43    PSNJ10544216-2104138
             JH imaging (to flux calibrate BG spectrum taken on previous night)

08.16    Switch to EFOSC2

08.23    SN2015D
             Focus
             2x1800s Gr#16 followup

09.40    LSQ15rw
             Gr#13 classification

10.00    LTT7379 spectrophotometric standard

             gr skyflats (aborted after 4x g' flats, as the sky was too bright).

Comments